OpenStudy (anonymous):

complete the solution of the equation. find the value of y when x equals -7. -7x-5y=54

OpenStudy (imer):

Just substitute "-7" in place of "x" and simplify \[-7(-7)-5y=54 \rightarrow y=-1\]

OpenStudy (anonymous):

thanks, so now would this work the same with: Complete the solution of the equation. Find the value of y when x=2. -3x-8=10

OpenStudy (imer):

yep, all you have to do is substitute "x" value in the equation and simplify. but I believe the equation you gave is incorrect, there is no "y" in the equation. Do you mean "y" instead of "8"?

OpenStudy (anonymous):

i forgot to fix it on the question but it should be 8y but i figured it out its y=-2